OpenStudy (anonymous):

-4sqrt(x+9)=20 can anyone help?

OpenStudy (anonymous):

\[-4 \sqrt{(x+9)} = 20\] Like this? At what part are you getting stuck?

OpenStudy (anonymous):

well i get x=16 but when i try to see if its true i get -20=20

OpenStudy (anonymous):

and yes like that

OpenStudy (anonymous):

Yeah, with that -4 outside the square root, I think it's undefined.

OpenStudy (anonymous):

yeah your probably right. extraneous solution?

OpenStudy (anonymous):

If we defined \(y\) as \(\sqrt{(x+9)}\), we would get \(-4y = 20\), with \(y = -5\), and there is no way \(\sqrt{(x+9)} = -5\)

OpenStudy (anonymous):

hmmm i dont know. but when we divide -4 we get \[\sqrt{x+9}=-5\]

OpenStudy (anonymous):

\(\large{\sqrt{-x} = i \sqrt{x}}\) right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Okay, can we use that somehow?

OpenStudy (anonymous):

Oh wait... We can square both sides, right?

OpenStudy (anonymous):

hmmm yeah

OpenStudy (anonymous):

and we get x+9=25 because -5x-5=25

OpenStudy (anonymous):

Right, so you were right with 16; it is an extraneous root.

OpenStudy (anonymous):

yeah it looks like that then and thanks man

OpenStudy (anonymous):

Hey, no problem. :)